Technical Staff Blog

Last update on .

network-2400px.png This was supposed to be a trouble free morning... just power up the switch we moved downstairs yesterday and add our redundancy back in. It seems to be a trend with this project, things not going quite as we planned...

07:14: The second distribution switch was powered on at 7am. Unfortuantely, something happened that caused our edge switches, the ones directly connecting departmental workstations, to enter an error state. Servers and the firewalls are unaffected by this outage, but desktops will be unable to reach the network until the issue is resolved. CIS is investigating.

07:15: Our secondary firewall is up and operational again. We are missing the management interface though, Max is investigating. This interface is not used for normal firewall traffic, so it should not have any impact on departmental users.

08:09: Edge network gear is starting to come back. CIS needs to physically connect to and log into every single switch to reset the uplinks, so it may be a bit before we are fully functional again.

08:23: CIS believes they have reset all the switches. We are digging our way through our monitoring system to confirm everything is back. In the mean time, if you are continuing to experience issues with network connetivity, please reach out to a Tstaff member (try out our "problem line" at 401-863-7629) or email problem@cs.brown.edu.